### v2

# Time-Varying Factor-Augmented Models for Volatility Forecasting

Abstract
Accurate volatility forecasts are vital in modern finance for risk management, portfolio allocation, and strategic decision-making. However, existing methods face key limitations. Fully multivariate models, while comprehensive, are computationally infeasible for realistic portfolios. Factor models, though efficient, primarily use static factor loadings, failing to capture evolving volatility co-movements when they are most critical. To address these limitations, we propose a novel, model-agnostic Factor-Augmented Volatility Forecast framework. Our approach employs a time-varying factor model to extract a compact set of dynamic, cross-sectional factors from realized volatilities with minimal computational cost. These factors are then integrated into both statistical and AI-based forecasting models, enabling a unified system that jointly models asset-specific dynamics and evolving market-wide co-movements. Our framework demonstrates strong performance across two prominent asset classes—large-cap U.S. technology equities and major cryptocurrencies—over both short-term (1-day) and medium-term (7-day) horizons. Using a suite of linear and non-linear AI-driven models, we consistently observe substantial improvements in predictive accuracy and economic value. Notably, a practical pairs-trading strategy built on our forecasts delivers superior risk-adjusted returns and profitability, particularly under adverse market conditions.
Keywords Volatility Forecasting, Factor Models, Long Short-Term Memory (LSTM), HAR, MIDAS, Volatility‑Based Pairs Trading
1 Introduction
From tick‑by‑tick hedging to long‑term investment strategy, reliable volatility forecasts are indispensable for guiding financial decisions. A vast literature—ranging from linear ARCH engle1982autoregressive , GARCH bollerslev1986generalized , and HAR corsi2009simple to nonlinear neural networks such as LSTMs hochreiter1997long —is typically applied on an asset-by-asset basis, modeling each asset in isolation. This univariate approach overlooks market-wide shocks that propagate rapidly across assets hamao1990correlations , often biasing forecasts and understating systemic risk. Although multivariate extensions such as BEKK‑GARCH engle1995multivariate , DCC‑GARCH engle2002dynamic , Conditional Autoregressive Wishart golosnoy2012conditional , and deep multivariate LSTM architectures capture shared dynamics, they are rarely used in real‑world settings due to parameter explosion, fragile estimation, and heavy tuning burdens bauwens2006multivariate ; engle2019large .
Factor models offer a natural remedy by efficiently summarizing co-movements through a low-dimensional set of latent drivers. Yet, two key limitations hinder their utility in volatility forecasting. First, most factor models are designed for returns, not volatility. Canonical linear frameworks such as Fama-French and Carhart fama1993common ; carhart1997persistence , and their modern machine‑learning successors gu2020empirical , all distill factors from returns data. Consequently, the rich patterns of commonality in volatility—which often exhibit their own distinct dynamics—remain largely unexplored. Second, although a few recent pioneering studies do extend factor models to realized volatility atak2013factor ; ding2025multiplicative , they still impose static factor loadings, which fail to capture time-varying interdependencies—an especially critical flaw given that volatility co-movement evolves over time del2008dynamic . Although recent econometric advances do provide time-varying loadings estimators mikkelsen2019consistent , these techniques have so far been confined exclusively to asset returns.
To bridge these gaps, we propose a model-agnostic Factor-Augmented Volatility Forecast framework that captures evolving market-wide volatility shocks with computational tractability. Unlike traditional factor approaches with static loadings, our method employs a time-varying factor model to extract latent factors directly from realized volatilities. This dynamic approach ensures that both the factors and their loadings adapt in real time to shifting market regimes. Compared to conventional multivariate techniques, our framework achieves this enhanced market awareness while preserving the parsimony of the base model, furthermore, introducing only a minimal number of extra parameters. The result is a robust, accurate, and interpretable framework that delivers computationally efficient volatility forecasts, suitable for practical deployment in dynamic financial environments.
We evaluate the framework on two representative markets—U.S. technology equities and cryptocurrencies—for both 1-day and 7-day forecast horizons. Across all model-market-horizon combinations, factor augmentation consistently improves performance. Out-of-sample coefficient of determination ($R^{2}$) increases by up to 12.8% for equities and 22.8% for cryptocurrencies, with corresponding reductions in Mean Squared Error (MSE) and Quasi-likelihood loss (QLIKE). Furthermore, a volatility-scaled pairs-trading backtest confirms the framework’s tangible economic value. Most strikingly, in a challenging market period, augmentation reversed an unprofitable strategy from an annualized loss of -5.5% into a +7.3% gain, flipping its Sharpe ratio from negative to positive. In brief, our key contributions are as follows:
  • We introduce the first volatility forecasting framework that systematically integrates dynamic, cross-sectional information through factor augmentation. Our approach combines three unique innovations: (i) direct factor extraction from realized volatilities, distilling interpretable factors that capture volatility-specific commonalities; (ii) time-varying loadings that adapt to evolving market interdependencies; and (iii) a computationally efficient augmentation mechanism that ensures both robustness and scalability.
  • Our framework achives superior statistical accuracy across all tested scenarios, and a practical pairs-trading backtest confirms that these accuracy gains translate directly into higher risk-adjusted returns, generating real-world economic value.
  • We provide a practical, end-to-end guide to factor extraction, selection, and integration, offering a clear roadmap for tailoring the framework to any forecasting model, horizon, or asset class.
2 Related Work
Research in discrete-time volatility forecasting has evolved significantly. However, existing models present common trade-offs, which our framework is designed to address.
Univariate Models. Frameworks from early AR models to more sophisticated ARCH engle1982autoregressive , GARCH bollerslev1986generalized ; bollerslev1987conditionally , and HAR corsi2009simple have long been pillars of the field due to their interpretability. Extensions have incorporated features like mixed-frequency data conrad2020two ; santos2014volatility and long-memory processes baillie1996long . However, by modeling each asset in isolation, these models suffer from a "univariate blind spot," ignoring the systemic co-movements that are crucial drivers of market-wide risk herskovic2016common ; hamao1990correlations .
Multivariate Models. Multivariate models are designed to capture asset interplay directly. Within the GARCH family, the BEKK specification engle1995multivariate guarantees positive-definite covariance matrices, while the Dynamic Conditional Correlation (DCC) model engle2002dynamic offers a more feasible estimation process. AR-style alternatives like the Conditional Autoregressive Wishart (CAW) golosnoy2012conditional provide further flexibility. While theoretically comprehensive, these models are often impractical for large systems due to the explosion of parameter count bauwens2006multivariate and numerically unstable estimation engle2019large .
Machine Learning Models. A parallel track has leveraged machine learning to capture the complex, nonlinear volatility patterns. Deep learning models like LSTM hochreiter1997long have shown superior performance over classical models by modeling long-range dependencies in financial time series liu2019novel ; mcnally2018predicting . However, these models are not immune to the "curse of dimensionality." When applied in a fully multivariate setting, they face excessive parameterization, high overfitting risk, and prohibitive computational costs ge2023review , inheriting the same fundamental limitations as their statistical counterparts.
Factor Models. Factor methods are the backbone of modern portfolio theory, distilling hundreds of correlated returns into a handful of common drivers and powering scalable asset allocation, hedging, and attribution. Widely used examples range from statistical factor‑GARCH variants such as OGARCH alexander2000primer and GO‑GARCH van2002go to recent factor‑structured deep networks feng2024deep . While these models achieve parsimony by using latent factors, most of the existing work targets returns gu2020empirical ; factor models for volatility remain comparatively sparse and always assume time‑invariant loadings atak2013factor ; ding2025multiplicative .
Our research directly confronts these trade-offs by proposing a factor augmentation framework. This approach enhances existing volatility models with a set of dynamic, cross-sectional factors, enabling the framework to capture systemic, time-varying market interdependence without adding additional computational overhead, thereby balancing predictive accuracy with practical feasibility.
3 Data and Volatility Measurement
3.1 Data
The empirical analysis is conducted on two high-frequency datasets, encompassing U.S. equities and major cryptocurrencies.
The equity dataset focuses on five prominent firms within the S&P 500 Information Technology Sector, selected for their large market capitalization and historical sectoral comovements: Microsoft (MSFT), Advanced Micro Devices (AMD), Intel (INTC), Oracle (ORCL), and Cisco (CSCO). For these firms, we obtained high-frequency Consolidated Trade and Quote records from Wharton Research Data Service, covering the period from January 2018 to December 2020 at millisecond precision.
The cryptocurrency dataset, sourced from Kaiko, covers January 2018 to April 2021. This dataset comprises detailed, high-frequency trading data for five leading cryptocurrencies: Bitcoin (BTC), Ethereum (ETH), Ripple (XRP), Cardano (ADA), and Litecoin (LTC). Kaiko compiles transaction-level records from major digital asset exchanges, providing comprehensive data including timestamps, trade volumes, and direction.
3.2 Volatility Measurement
3.2.1 Daily Realized Volatility
Following andersen2001distribution , we estimate daily realized volatility (RV) nonparametrically from high-frequency midpoint prices. After filtering out non-positive, crossed, and spurious quotes, midpoint prices for asset $i$ on day $t$ are sampled every five minutes:
$\text{MID}_{i,t}=\frac{\text{BID}_{i,t}+\text{ASK}_{i,t}}{2}.$ (1)
The daily RV is then calculated as the square root of the sum of squared intraday returns:
$RV_{i,t}=\sqrt{\sum_{s=1}^{N}\left[\ln\left(\frac{\text{MID}_{i,t_{s}}}{\text{MID}_{i,t_{s-1}}}\right)\right]^{2}}\;,$ (2)
where $\mathrm{MID}_{i,t_{s}}$ is the mid‐price at the $s$th five‐minute interval, $N=78$ for a standard equity trading day and $N=288$ for 24‐hour cryptocurrency markets.
3.2.2 Aggregated Realized Volatility
To capture medium-term risk dynamics, we construct weekly realized volatility by averaging daily RVs over a 7-day window ($h=7$) following (corsi2009simple, ; bollerslev2018risk, ):
$RV_{t}^{7}=\frac{1}{7}\sum_{j=0}^{6}RV_{t-j}\;.$ (3)
This 7‑day moving average smooths daily noise, yielding a stable volatility gauge that matches common weekly rebalancing and reporting cycles.
4 Methodology
4.1 Volatility Factor Analysis Across Assets
4.1.1 Time-Varying Factor Model
Let $\mathbf{y}_{t}=(RV_{1,t},\dots,RV_{p,t})$ denote the $p$-dimensional vector of realized volatilities at time $t=1,\ldots,T$. Following the time‑varying matrix factor model of chen2024timevaryingmatrixfactormodels , we posit the locally‑smooth factor representation:
$\mathbf{y}_{t}\;=\;\mathbf{\Lambda}_{t}\mathbf{f}_{t}+\mathbf{\varepsilon}_{t},$ (4)
where $\mathbf{f}_{t}\in\mathbb{R}^{k}$ is the $k$-dimensional latent factors, and $\mathbf{\varepsilon}_{t}\in\mathbb{R}^{p}$ is the idiosyncratic errors, assumed to be uncorrelated with $\mathbf{f}_{t}$ but allowed to exhibit weak serial dependence. To accommodate a wide range of potential smooth temporal variation, we specify $\mathbf{\Lambda}_{t}\in\mathbb{R}^{p\times k}$ to be the time‑varying loading matrix. Specifically, we model $\mathbf{\Lambda}_{t,i}=\mathbf{\Lambda}_{i\cdot}(t/T)$, where $\mathbf{\Lambda}_{i\cdot}(\cdot)$ is an unknown smooth function of $t/T$ on $[0,1]$ for each $i\in[p]$.
Let $\mathbf{\Sigma}_{f}=\mathbb{E}(\mathbf{f}_{t}\mathbf{f}_{t}^{\top})$ and $\mathbf{\Sigma}_{\varepsilon}=\mathbb{E}(\varepsilon_{t}\varepsilon_{t}^{\top})$. The second uncentered moment satisfies
$\mathbf{\Sigma}_{y,t}:=\mathbb{E}(\mathbf{y}_{t}\mathbf{y}_{t}^{\top})=\mathbf{\Lambda}_{t}\mathbf{\Sigma}_{f}\mathbf{\Lambda}_{t}^{\top}+\mathbf{\Sigma}_{\varepsilon}.$ (5)
Under mild regularity conditions, the rank‑$k$ signal component $\mathbf{\Lambda}_{t}\mathbf{\Sigma}_{f}\mathbf{\Lambda}_{t}^{\top}$ implies that the top $k$ eigenvectors of $\mathbf{\Sigma}_{y,t}$ span the column space of $\mathbf{\Lambda}_{t}$ up to rotation, as established in chen2024timevaryingmatrixfactormodels .
In principle, one could estimate $\mathbf{\Sigma}_{y,t}$ using kernel smoothing in $t/T$ as in chen2024timevaryingmatrixfactormodels . For practical implementation, we instead adopt a rolling‑window (uniform‑kernel) approximation. Fix a window length $n$ and define $W_{t}=\{t-n+1,\dots,t\}$. We estimate the local covariance matrix by
$\widehat{\mathbf{\Sigma}}_{y,t}=\frac{1}{n}\sum_{s\in W_{t}}\mathbf{y}_{s}\mathbf{y}_{s}^{\top}=\sum_{s=1}^{T}K_{n}(s,t)\,\mathbf{y}_{s}\mathbf{y}_{s}^{\top},\;K_{n}(s,t)=\frac{1}{n}\,\mathbf{1}\{s\in W_{t}\}.$
We extract the top-k eigenvectors $\hat{\mathbf{u}}_{1,t},\ldots,\hat{\mathbf{u}}_{k,t}$ of $\widehat{\mathbf{\Sigma}}_{y,t}$. The estimated loading space and the contemporaneous factors are then defined by
$\widehat{\mathbf{\Lambda}}_{t}\;=\;\sqrt{p}\,[\hat{\mathbf{u}}_{1,t},\ldots,\hat{\mathbf{u}}_{k,t}],\qquad\widehat{\mathbf{f}}_{t}\;=\;\frac{1}{p}\,\widehat{\mathbf{\Lambda}}_{t}^{\top}\mathbf{y}_{t}.$
We apply this time-varying factor model separately to equities and cryptocurrencies, producing two distinct sets of $\widehat{\mathbf{\Lambda}}_{t}\text{ and }\widehat{\mathbf{f}}_{t}$.
4.1.2 Divergent Factor Structures: Concentration in Crypto vs. Diffusion in Equity
Both the equity and cryptocurrency markets exhibit hierarchical factor structures, yet they diverge significantly in risk concentration. This structural difference enables clear economic interpretations of the factors and provides a direct rationale for our horizon-dependent factor selection.
The equity market’s risk is diffuse. Its leading factor, $\hat{\mathbf{f}}_{t}^{(1)}$, represents broad technology sector risk, explaining 45.2% of the total variance, with firms like Oracle loading most heavily at 0.96. However, as the gradual decay in the scree plot Figure 1(b) illustrates, a single factor is insufficient. Subsequent components, from $\hat{\mathbf{f}}_{t}^{(2)}\text{ to }\hat{\mathbf{f}}_{t}^{(5)}$, capture firm-specific risks for companies like Microsoft and Intel, meaning a broader set of factors is needed to comprehensively describe market dynamics.
In stark contrast, cryptocurrency’s risk is highly concentrated. A single, dominant systemic factor $\hat{\mathbf{f}}_{t}^{(1)}$ serves as the primary market-wide driver, accounting for a commanding 79.1% of total variance. This factor is closely linked to the Bitcoin cycle, with its overwhelming influence confirmed by the sharp "elbow" in Figure 1(d). While subsequent factors isolate the dynamics of major altcoins like Ethereum and Ripple, their marginal contributions are more modest, highlighting a market driven by one principal force.
These distinct factor structures provide a clear rationale for our horizon-dependent factor selection. For short-term (1-day) forecasts, we select only the single, dominant factor for each asset class to maximize the signal-to-noise ratio. For medium-term (7-day) forecasts, we employ an explained-variance threshold jolliffe2016principal ; pukthuanthong2009global . This involves selecting the minimum number of factors required to meet a predetermined cumulative variance target, which we tailor to each market’s unique structure (e.g., 85% for equities and 90% for crypto). This data-driven approach ensures we capture significant secondary dynamics without overfitting on noise.
(a) Factor Loadings for Equity
(b) Explained Variance for Equity
(c) Factor Loadings for Crypto
(d) Explained Variance for Crypto
Figure 1: Factor Loadings and Scree Plots for Equities and Crypto
4.2 Factor-Augmented Forecasting Models
To demonstrate the flexibility and model-agnostic nature of our framework, we apply it uniformly to four benchmark models. The models forecast a common target variable, $Y_{t+h}$, across two horizons $h\in\{1,7\}$: the next day’s daily volatility $Y_{t+1}=RV_{t+1}$ and the aggregated weekly volatility for the upcoming week $Y_{t+7}=RV^{7}_{t+7}$.
The forecasts are based on two estimation approaches: an expanding window for the three statistical models and a fixed 80:20 train-test split for the more computationally intensive LSTM.
4.2.1 Autoregressive (AR)
The AR model regress the target volatility $Y_{t+h}$ on the five most recent daily RVs to capture short-memory dynamics over a full trading week without over-parameterization:
$Y_{t+h}=\beta_{0}+\sum_{i=0}^{4}\beta_{i}RV_{t-i}+\epsilon_{t}.$ (6)
Given the AR model’s simple linear structure, the factor-augmented version extends the baseline by directly incorporating the top-$S$ factors as additional predictors:
$Y_{t+h}\;=\;\beta_{0}+\sum_{i=0}^{4}\beta_{i}RV_{t-i}\;+\;\sum_{j\in[S]}\gamma_{j}\,\widehat{\mathbf{f}}_{j,t}\;+\;\epsilon_{t}.$ (7)
4.2.2 Heterogeneous Autoregressive (HAR)
The HAR model captures multi-scale volatility persistence using daily, weekly, and monthly components corsi2009simple :
$Y_{t+h}=\beta_{0}+\beta^{d}RV_{t}+\beta^{w}RV^{7}_{t}+\beta^{m}RV^{30}_{t}+\epsilon_{t},$ (8)
where $RV^{d}_{t}$, $RV^{w}_{t}$, and $RV^{m}_{t}$ represent daily, weekly, and monthly volatilities, respectively.
To align with the HAR’s multi-scale structure, our augmentation introduces factors constructed over corresponding daily and weekly horizons:
$\begin{split}Y_{t+h}\;=&\;\beta_{0}+\beta^{d}RV_{t}+\beta^{w}RV^{7}_{t}+\beta^{m}RV^{30}_{t}\\ &\;+\;\sum_{j\in[{S}^{d}]}\gamma^{d}_{j}\,\widehat{\mathbf{f}}^{\,d}_{j,t}\;+\;\sum_{r\in[{S}^{w}]}\gamma^{w}_{r}\,\widehat{\mathbf{f}}^{\,w}_{r,t}\;+\;\epsilon_{t},\end{split}$ (9)
where $\widehat{\mathbf{f}}^{d}_{j,t}$ and $\widehat{\mathbf{f}}^{w}_{r,t}$ denote the top-$S^{d}$ daily factors and the top-$S^{w}$ weekly factors, respectively.
4.2.3 Mixed Data Sampling (MIDAS)
The MIDAS model captures long-memory dynamics by incorporating many past observations through a parsimonious weighting scheme, specified as a Beta-polynomial distributed lag ghysels2007midas :
$Y_{t+h}=\beta_{0}+\beta_{1}\left[a(1)^{-1}a(L)RV_{t}\right]+\epsilon_{t},$ (10)
where lag weights $a(L)$ follow:
$a_{i}=\left(\frac{i}{k}\right)^{\theta_{1}-1}\left(1-\frac{i}{k}\right)^{\theta_{2}-1}\frac{\Gamma(\theta_{1}+\theta_{2})}{\Gamma(\theta_{1})\Gamma(\theta_{2})},\quad i=1,\ldots,k,$ (11)
with normalization $a(1)=\sum_{i=1}^{k}a_{i}$. $\theta_{1}=1$ and $\theta_{2}$ are optimized via in-sample grid search. Default $k=30$ is extended with $k\in\{30,50,80\}$ for longer horizons.
To align with the MIDAS’s core principle of parsimoniously weighting long-lag structures, we augment each factor through its own analogous Beta-polynomial $b(L)$:
$Y_{t+h}\;=\;\beta_{0}+\beta_{1}\!\bigl[a(1)^{-1}a(L)RV_{t}\bigr]+\sum_{j\in[S]}\gamma_{j}\bigl[b(1)^{-1}b(L)\,\widehat{\mathbf{f}}_{j,t}\bigr]+\epsilon_{t}.$ (12)
4.2.4 Long Short-Term Memory (LSTM)
Long Short‑Term Memory network is a gated recurrent architecture designed to capture long‑range dependencies in time‑series data. Each input vector concatenates a 7‑day window of daily realized volatility with the contemporaneous factor. This sequence passes through a stack of three LSTM layers, which learn the joint temporal patterns of volatility and factors. The final hidden state feeds a dense output layer that forecasts the target volatility $Y_{t+h}$.
4.3 Evaluation Metrics
Model performance is examined through three lenses: (i) statistical accuracy, (ii) statistical significance, and (iii) economic value.
4.3.1 Prediction Accuracy
Prediction accuracy is assessed with three widely used metrics: $R^{2}$, MSE, and QLIKE, which respectively capture explained variance, average prediction error, and distributional accuracy respectively patton2011volatility ; zhang2024volatility :
$R^{2}$ $=1-\sum_{t}\left(\frac{RV^{h}_{t+h}-\widehat{RV}^{h}_{t+h}}{RV^{h}_{t+h}-\overline{RV}^{h}_{t+h}}\right)^{2},$ (13)
$\mathrm{MSE}$ $=\frac{1}{T}\sum_{t}\left(RV^{h}_{t+h}-\widehat{RV}^{h}_{t+h}\right)^{2},$ (14)
$\mathrm{\mathrm{QLIKE}}$ $=\frac{1}{T}\sum_{t}\!\left[\frac{RV^{h}_{t+h}}{\widehat{RV}^{h}_{t+h}}-\ln\!\left(\frac{RV^{h}_{t+h}}{\widehat{RV}^{h}_{t+h}}\right)-1\right],$ (15)
4.3.2 Economic Implication
We assess practical relevance using the out-of-sample Utility-of-Wealth ($\mathrm{UoW}$) metric:
$\widehat{\mathrm{UoW}}=\frac{1}{T}\sum_{t=1}^{T}\!\left[\frac{SR^{2}}{\gamma}\frac{RV^{7}_{t+7}}{\widehat{RV}^{7}_{t+7}}-\frac{SR^{2}}{2\gamma}\!\left(\frac{RV^{7}_{t+7}}{\widehat{RV}^{7}_{t+7}}\right)^{2}\right],$ (16)
where we follow bollerslev2018risk in setting the target Sharpe ratio to $\mathrm{SR}=0.4$, a typical long-run estimate for volatility-timing strategies, and choose risk aversion $\gamma=2$ to represent moderate investor preferences. Higher $\mathrm{UoW}$ denotes greater expected utility for a volatility-timing investor.
4.3.3 Statistical Significance
We assess systematic improvements with the Diebold–Mariano testdiebold2002comparing . For each horizon $h$, we compute the one‐period loss differential:
$d_{t}\;=\;L\bigl(RV^{h}_{t+h},\widehat{RV}^{h}_{t+h,\mathrm{aug}}\bigr)\;-\;L\bigl(RV^{h}_{t+h},\widehat{RV}^{h}_{t+h,\mathrm{RW}}\bigr),$ (17)
where
[…]
The $\mathrm{DM}$ statistic:
$\mathrm{DM}=\frac{\bar{d}}{\sqrt{\widehat{\mathrm{Var}}(\bar{d})}},\quad\bar{d}=\frac{1}{T}\sum_{t=1}^{T}d_{t},$ (19)
is asymptotically $\mathcal{N}(0,1)$ under $H_{0}:E[d_{t}]=0$. A positive DM value indicates that the factor-augmented forecasts significantly outperform the unaugmented benchmarks.
5 Experiment Result
Table 1: Cryptocurrency Volatility Forecast Accuracy: 1- and 7-Day Out-of-Sample Results
Panel A: Crypto in 1‐Day Horizon
BTC ETH XRP ADA LTC
$\bm{R^{2}}$ MSE QLIKE $\bm{R^{2}}$ MSE QLIKE $\bm{R^{2}}$ MSE QLIKE $\bm{R^{2}}$ MSE QLIKE $\bm{R^{2}}$ MSE QLIKE
AR 43.60 3.27 0.0802 40.34 3.50 0.0687 45.06 7.69 0.0842 41.33 4.16 0.0499 39.95 4.13 0.0658
AR‐Aug 47.57 2.99 0.0770 42.34 3.38 0.0679 48.58 7.15 0.0789 42.01 4.13 0.0498 40.02 4.12 0.0654
HAR 39.52 3.12 0.0831 35.93 3.41 0.0691 47.58 6.98 0.0779 40.51 4.16 0.0497 34.08 4.09 0.0640
HAR‐Aug 43.93 0.88 0.0797 40.43 3.19 0.0681 49.73 6.63 0.0778 41.09 4.14 0.0497 37.41 3.85 0.0640
MIDAS(k=30) 42.20 2.96 0.0828 39.07 3.29 0.0691 45.95 7.21 0.0843 40.94 4.19 0.0508 33.58 4.03 0.0683
MIDAS‐Aug(k=30) 42.32 2.96 0.0827 38.93 3.30 0.0693 46.00 7.21 0.0844 40.94 4.19 0.0508 34.12 4.01 0.0679
LSTM 53.88 2.13 0.0603 51.24 2.75 0.0551 50.73 13.20 0.0824 45.12 6.12 0.0550 51.64 3.66 0.0530
LSTM‐Aug 57.70 1.99 0.0564 55.72 2.64 0.0506 53.57 12.66 0.0800 47.34 6.08 0.0529 53.64 3.65 0.0519
Panel B: Crypto in 7‐Day Horizon
AR 48.63 1.87 0.0677 39.98 2.08 0.0552 41.78 5.53 0.0801 38.37 2.89 0.0426 39.21 2.24 0.0475
AR‐Aug 50.89 1.78 0.0648 41.85 2.03 0.0544 43.07 5.49 0.0792 40.70 2.78 0.0414 42.38 2.14 0.0453
HAR 44.96 2.00 0.0793 38.25 2.15 0.0571 45.63 5.24 0.0782 36.71 2.96 0.0442 39.14 2.26 0.0492
HAR‐Aug 47.38 1.94 0.0737 37.90 2.17 0.0577 46.58 5.29 0.0782 38.03 2.90 0.0435 40.36 2.22 0.0474
MIDAS(k=30) 49.40 1.84 0.0670 40.94 2.06 0.0546 43.26 5.47 0.0796 38.92 2.86 0.0424 40.82 2.20 0.0469
MIDAS‐Aug(k=30) 53.26 1.70 0.0634 44.04 1.95 0.0516 47.23 5.09 0.0724 47.01 2.47 0.0390 45.08 2.07 0.0438
MIDAS(k=50) 49.64 1.83 0.0670 40.90 2.06 0.0546 43.26 5.52 0.0804 39.00 2.86 0.0423 40.69 2.20 0.0470
MIDAS‐Aug(k=50) 52.95 1.71 0.0621 43.80 1.95 0.0514 47.34 5.05 0.0727 47.29 2.47 0.0389 44.30 2.07 0.0443
MIDAS(k=80) 49.58 1.83 0.0673 40.96 1.95 0.0516 42.73 5.09 0.0724 38.92 2.48 0.0390 40.80 2.04 0.0438
MIDAS‐Aug(k=80) 52.95 1.71 0.0628 43.72 1.96 0.0518 47.33 5.08 0.0736 47.79 2.45 0.0384 44.31 2.07 0.0447
LSTM 59.31 1.37 0.0419 46.25 2.28 0.0441 42.81 1.43 0.1222 37.32 6.28 0.0629 52.39 2.38 0.0401
LSTM‐Aug 61.66 1.30 0.0374 51.89 2.03 0.0385 44.25 1.41 0.1186 41.16 6.19 0.0616 59.83 2.38 0.0340
Note: All MSE values are in units of $10^{-4}$.   Improvement:  $\geq+15\%$  $\geq+10\%$  $\geq+5\%$  $0\%$  $\leq-5\%$
5.1 A Universal Performance Boost Across All Models and Horizons
Factor augmentation universally and significantly boosts forecast accuracy across every model, asset class, and horizon tested, proving its effectiveness for both short- and medium-term predictions, as shown in Tables 1 and 4.
Specifically, in 1‑day forecasts, the benefits of factor augmentation are most pronounced for short-memory models, with HAR and LSTM boosting equity $R^{2}$ by up to 10% and crypto $R^{2}$ by a commanding 9–12%. The MIDAS model, as expected from its long-memory design, shows more modest improvement.
At the 7-day horizon, long‑memory models exhibit more substantial gains. While AR and HAR still show solid improvements, the augmented LSTM achieves an accuracy boost of up to 14%. The MIDAS model now stands out, delivering remarkable gains: its $R^{2}$ in equities surges by 11–13%, and by as much as 23% in cryptocurrencies, with corresponding reductions in MSE and QLIKE.
These results demonstrate the versatile power of factor augmentation, which consistently improves forecast accuracy, especially when a model’s architecture aligns with the task horizon.
5.2 Value Across Markets: Robust in Equities, Super‑Charged in Crypto
Factor augmentation framework proves its power and versatility across fundamentally different market structures, delivering robust gains in the diffuse equity market and a super-charged performance boost in the highly concentrated crypto market.
In the challenging equity market, where volatility drivers are highly dispersed, our framework delivers substantial value. For instance, as shown in Table 4, the augmented 7-day MIDAS model boosts its $R^{2}$ by an average of 4.7%, while the 1-day LSTM improves its accuracy by 4.1%, proving the framework’s reliability even in such diffuse regimes.
In cryptocurrency markets, the framework’s competitive edge is even sharper. Here, where the top three factors capture nearly 95% of market variance, the performance gains are dramatic: as detailed in Table 1, the 7‑day MIDAS model’s $R^{2}$ soars by an average of 11.1% and the 1-day HAR model’s by 7.9%—both more than doubling the improvements seen in equities.
These results confirm that factor augmentation is not only a reliable enhancer in different markets, but also becomes especially powerful when risk drivers are concentrated.
5.3 Statistically and Economically Significant Gains
Factor augmentation delivers gains that are both substantial and statistically significant in both predictive accuracy and economic value. We demonstrate this on the crypto 7-day horizon task, where our framework shows the largest improvements. Three unaugmented baselines are employed: the Random Walk, an "Individual" model with asset-specific coefficients, and an "Panel" model that share coefficients across assets.
The framework’s edge in predictive accuracy is both substantial and statistically unequivocal. As shown in Table 5(a), while the augmented models’ consistently higher $R^{2}$ offer strong initial evidence, the positive statistics from an MSE-based DM test provide formal proof. This confirms that the forecasting improvements are not due to random chance but are a robust gain of our methodology.
Crucially, these statistical superiority translate directly into significant economic value. As detailed in Table 5(b), the augmented models consistently deliver unambiguously higher UoW outcomes, and a corresponding UoW-based DM test formally validates that these gains are statistically significant, providing clear evidence of the framework’s real-world benefits for investors.
6 Portfolio Backtest
To rigorously assess the framework’s financial value, we conduct a volatility‑scaled pair‑trading backtest.
6.1 Cointegration Test
To identify a stationary, mean-reverting spread suitable for our pairs-trading strategy, we perform a two-stage cointegration analysis on the price series within each market.
We apply the Johansen trace test to each five-asset group. While it finds no evidence of cointegration at any rank for the equity group, the crypto sector indicates a cointegration rank of $r=2$.
Table 2: Cointegrating Vector Coefficients for Crypto
ADA BTC ETH LTC XRP
CEV1 +2.62+2.62 +3.69+3.69 4.60-4.60 0.99-0.99 +2.23+2.23
CEV2 +5.15+5.15 0.19-0.19 5.64-5.64 +1.82+1.82 3.51-3.51
Table 2 reports the first two cointegrating vectors ($\text{CEV}_{1,2}$), both of which load ADA and ETH with opposite signs, indicating that the dominant mean-reverting combination is the ADA–ETH spread. To further validate this pairing, we conduct pairwise Engle–Granger tests on each residual and find only ADA–ETH stationary at the 5% level ($p=0.045$). Consequently, our backtest focuses exclusively on the ADA–ETH spread, excluding all other pairs.
6.2 Volatility‑Based Pairs Trading
Our backtest framework simulates a pairs trading strategy on the cointegrated spread between ADA-ETH. The log-price spread $S_{t}$ is defined as:
$S_{t}=\ln P_{\text{ADA},t}-\beta\,\ln P_{\text{ETH},t},$ (20)
where the hedge ratio $\beta$ is estimated via rolling OLS regression. The performance is evaluated for all four benchmark models and is measured by annualized return and Sharpe ratio.
The performance is evaluated by final equity, cumulative return, and annualized Sharpe ratio.
6.2.1 Trading Signal Generation
Trading signals are generated from a rolling z-score of the spread, where $Z_{t}=\frac{S_{t}-\mu_{S_{t}}}{\sigma_{S_{t}}}$. We initiate a long position when $Z_{t}$ < -1.5 and a short position when $Z_{t}$ > +1.5, closing the position when the z-score reverts to zero. To limit turnover, each trade is held for at least one day. The rolling window length ($W$) is set to 70 days for the statistical models to balance responsiveness to regime changes with statistical stability, while a shorter window of 30 days is used for the LSTM to align with its input structure.
6.2.2 Volatility-based Position Sizing
Our strategy employs a volatility-targeting approach, where daily position sizes are scaled to achieve a 25% annualized volatility target. Each day, we estimate the spread volatility using our model’s forecasts for the component assets’ volatility $\widehat{\sigma}_{\text{ADA},t}$, $\widehat{\sigma}_{\text{ETH},t}$ and their return covariance:
$\widehat{\sigma}_{\text{spread},t}\;=\;\sqrt{252}\cdot\sqrt{\widehat{\sigma}_{\text{ADA},t}^{2}+\beta^{2}\widehat{\sigma}_{\mathrm{ETH},t}^{2}-2\beta\,\widehat{\mathrm{Cov}}(r_{\text{ADA},t},r_{\text{ETH},t})}.$ (21)
The position size is then set based on this forecast, subject to a 5:1 maximum leverage. The backtest is initialized with an equity of \$50,000 and incorporates a 5 basis point transaction cost per leg.
6.3 Backtest Results
To ensure a fair comparison, we evaluate statistical models on the full sample with an expanding window and the LSTM on a fixed 80:20 split. Each approach is benchmarked against a corresponding Random Walk baseline, with full results presented in Table 3.
Table 3: Out-of-Sample Backtest for ADA-ETH Pair
Panel A: Expanding Window Models
Model Metric Unaugmented Factor-Augmented Improvement
Random Walk (Full Sample) Portfolio Value (\$) 59,161.34
Ann. Return 5.32%
Ann. Sharpe Ratio 1.402
AR Portfolio Value 58,711.22 60,898.40 3.73%
Ann. Return 5.35% 6.61% 23.55%
Ann. Sharpe Ratio 1.437 1.443 0.42%
HAR Portfolio Value 56,311.89 57,298.08 1.75%
Ann. Return 3.86% 4.52% 17.10%
Ann. Sharpe Ratio 1.457 1.475 1.24%
MIDAS (k=30) Portfolio Value 58,339.32 58,121.55 -0.37%
Ann. Return 5.31% 5.04% -5.08%
Ann. Sharpe Ratio 1.460 1.485 1.71%
Panel B: Late Sample Models
Random Walk (Last 20% Window) Portfolio Value 49,125.74
Ann. Return -2.67%
Ann. Sharpe Ratio -0.18
LSTM Portfolio Value 47,427.84 53,306.58 12.40%
Ann. Return -5.48% 7.27% 232.66%
Ann. Sharpe Ratio -0.467 0.787 268.52%
6.3.1 Dual Gains: Amplifying Return and Sharpening Risk
Factor augmentation decisively boosts both profitability and risk-adjusted performance across all statistical models we examine.
The short-memory models realize the most impressive gains in raw profitability. Relative to its unaugmented version, the AR specification boosts its annualized return by a remarkable 23.6%. The HAR model shows a parallel improvement, with its return climbing by 17.1% and its Sharpe ratio rising from 1.457 to 1.475.
The long-memory MIDAS model, in contrast, demonstrates its value through a different channel: superior risk calibration. Although its raw return sees only modest gains—a result of its long-term design—the augmented strategy’s precise calibration of volatility drives its Sharpe ratio from 1.460 to a higher 1.485, signaling tighter risk management.
Collectively, this evidence confirms that factor augmentation is not merely an incremental tweak but a fundamental advance in volatility forecasting. It consistently amplifies reward, sharpens risk-adjusted performance, and sets a new benchmark for the field.
6.3.2 Robustness in All Market Conditions: Enhancing Gains, Reversing Losses.
Our framework proves its value in all market conditions, systematically enhancing gains in favorable periods while acting as a powerful shield during downturns.
During the full‐sample “bull‑market” window, when even a naïve RW was profitable, factor augmentation still delivered a distinct competitive edge. Every statistical model, when augmented, posted a higher Sharpe ratio, sharpening performance even when market conditions are already favorable.
The framework’s power becomes even more pronounced in adversity. During the challenging late-sample period, when the market turned negative and the RW benchmark lost 2.7% annually, the unaugmented LSTM suffered a -5.5% annualized loss, likely overfitting on patterns that no longer held. In a stunning reversal, our factor-augmented LSTM transformed this failing strategy into a resounding success, flipping the annualized loss into a +7.3% gain. This phenomenal turnaround propelled its Sharpe ratio from a negative -0.467 to a positive 0.787, unequivocally converting a losing proposition into a clear winner.
This remarkable dual capability—to amplify gains in good times and decisively reverse losses in bad—undeniably proves our framework’s practical utility for navigating the complexities of any market conditions.
7 Conclusion
This paper proposes a novel, model-agnostic Factor-Augmented Volatility Forecast framework, resolving the trade-off between computationally infeasible multivariate models, overly simplistic univariate models, and rigid static-factor models. By extracting a compact set of dynamic factors and their time-varying loadings directly from realized volatilities, our framework allows any baseline model to efficiently capture evolving market-wide co-movements. The framework’s robust performance—evidenced by substantial gains in forecast accuracy and superior risk-adjusted returns—combined with its computational efficiency and interpretability, establishes it as a valuable and practical solution for dynamic risk management.
Appendix
8 Equity Market: Forecast Performance
Table 4: Equity Volatility Forecast Accuracy: 1- and 7-Day Out-of-Sample Results
Panel A: Equity in 1‐Day Horizon
Models MSFT AMD INTC ORCL CSCO
$\bm{R^{2}}$ MSE QLIKE $\bm{R^{2}}$ MSE QLIKE $\bm{R^{2}}$ MSE QLIKE $\bm{R^{2}}$ MSE QLIKE $\bm{R^{2}}$ MSE QLIKE
AR 33.75 7.30 0.1756 32.07 8.79 0.0570 33.64 13.79 0.1115 36.49 23.47 0.1486 34.58 12.72 0.1133
AR‐Aug 34.26 7.30 0.1741 32.33 8.78 0.0566 34.01 13.57 0.1111 37.69 22.78 0.1443 34.78 12.55 0.1132
HAR 33.48 7.40 0.1740 32.87 8.72 0.0540 31.37 14.28 0.1139 37.11 23.37 0.1469 34.34 12.75 0.1087
HAR‐Aug 33.78 7.37 0.1755 32.52 8.84 0.0548 31.96 14.20 0.1135 37.21 23.21 0.1453 37.79 12.08 0.1049
MIDAS (k=30) 35.58 7.16 0.1759 34.58 8.61 0.0535 33.89 14.04 0.1134 38.19 23.39 0.1504 36.84 12.53 0.1083
MIDAS‐Aug 35.44 7.17 0.1771 34.49 8.62 0.0533 34.29 13.95 0.1129 38.60 23.46 0.1509 36.99 12.52 0.1083
LSTM 38.87 4.47 0.1237 27.51 13.22 0.0577 42.92 21.12 0.1103 27.10 54.25 0.1887 46.74 18.00 0.0887
LSTM‐Aug 39.33 4.43 0.1223 28.44 13.05 0.0564 44.16 20.73 0.1089 29.84 52.21 0.1876 48.08 17.79 0.0882
Panel B: Equity in 7‐Day Horizon
AR 44.99 3.48 0.0946 47.20 4.12 0.0269 44.03 6.71 0.0499 45.67 12.87 0.0730 47.85 6.12 0.0514
AR‐Aug 45.34 3.46 0.0964 47.43 4.10 0.0268 43.87 6.73 0.0503 46.81 12.59 0.0718 49.38 5.94 0.0504
HAR 44.39 3.86 0.1084 42.88 4.48 0.0257 37.83 8.36 0.0540 20.99 21.98 0.1154 40.32 8.14 0.0558
HAR‐Aug 45.71 3.81 0.1096 43.24 4.39 0.0258 40.07 8.06 0.0542 21.57 21.82 0.1036 45.27 7.47 0.0555
MIDAS (k=30) 47.26 3.33 0.0875 48.09 3.73 0.0254 45.53 6.21 0.0477 47.82 11.46 0.0741 50.02 5.61 0.0515
MIDAS‐Aug 49.61 3.18 0.0853 48.76 3.76 0.0251 46.29 6.12 0.0472 48.37 11.33 0.0711 55.53 4.99 0.0493
MIDAS(k=50) 46.35 3.41 0.0927 50.17 3.65 0.0251 45.36 6.27 0.0500 48.67 11.35 0.0842 55.30 5.05 0.0572
MIDAS‐Aug(k=50) 48.38 3.28 0.0859 50.49 3.63 0.0247 46.52 6.13 0.0472 48.70 11.35 0.0793 55.92 4.98 0.0545
MIDAS(k=80) 49.05 3.15 0.0885 50.85 3.76 0.0249 45.13 6.42 0.0496 48.22 11.95 0.0750 50.07 5.72 0.0526
MIDAS‐Aug(k=80) 51.36 3.01 0.0866 51.49 3.71 0.0239 45.68 6.35 0.0485 49.86 11.87 0.0728 56.50 4.98 0.0503
LSTM 46.38 3.55 0.1002 43.82 7.02 0.0337 42.98 15.13 0.0779 42.14 39.23 0.1491 49.47 13.01 0.0644
LSTM‐Aug 47.36 3.51 0.0964 44.82 6.91 0.0326 45.15 14.54 0.0766 42.74 33.29 0.1416 50.55 12.74 0.0628
Note: All MSE values are in units of $10^{-4}$.   Improvement:  $\geq+15\%$  $\geq+10\%$  $\geq+5\%$  $0\%$  $\leq-5\%$
9 Statistical Significance
(a) Out-of-Sample $R^{2}$ and MSE-based DM Test for 7-day Crypto
Asset Model Random Walk AR5 HAR MIDAS (k=30) MIDAS (k=50) MIDAS (k=80) LSTM
Panel A: Out-of-Sample $\bm{R^{2}}$ (in %)
BTC Individual 32.48 48.63 44.96 49.40 49.64 49.58 59.31
Panel 48.02 47.65 48.52 47.59 45.62 57.91
Factor-Augmented 50.89 47.38 53.26 52.95 52.95 61.66
ETH Individual 17.27 39.98 38.25 40.94 40.90 40.96 46.25
Panel 40.83 39.54 41.06 41.13 39.37 48.65
Factor-Augmented 41.85 37.90 44.04 43.80 43.72 51.89
XRP Individual 22.94 41.78 45.63 43.26 43.26 42.73 42.81
Panel 42.68 46.76 42.86 42.27 41.10 46.81
Factor-Augmented 43.07 46.58 47.23 47.34 47.33 44.25
ADA Individual 12.32 38.37 36.71 38.92 39.00 38.92 37.32
Panel 40.15 42.36 40.51 40.76 39.48 48.96
Factor-Augmented 40.70 38.03 47.01 47.29 47.79 41.16
LTC Individual 20.23 39.21 39.14 40.82 40.69 40.80 52.39
Panel 39.37 41.22 39.84 38.87 36.58 56.54
Factor-Augmented 42.38 40.36 45.08 44.30 44.31 59.83
Panel B: MSE-based Diebold–Mariano Test Statistics
BTC Aug vs. RW NA 2.46 1.25 3.34 3.36 3.32 3.71
Aug vs. Ind 3.13 1.29 2.18 2.21 2.06 4.32
ETH Aug vs. RW NA 3.07 2.34 4.60 4.56 4.48 3.56
Aug vs. Ind 1.65 1.61 2.10 2.19 1.95 2.71
XRP Aug vs. RW NA 2.60 3.09 3.50 3.57 3.55 2.94
Aug vs. Ind 1.89 1.78 1.79 1.77 1.68 2.13
ADA Aug vs. RW NA 3.29 3.35 3.80 3.76 3.84 1.95
Aug vs. Ind 2.14 1.14 2.11 1.79 1.70 1.61
LTC Aug vs. RW NA 2.18 1.94 3.78 3.78 3.79 3.12
Aug vs. Ind 3.94 1.52 1.27 1.10 1.27 2.09
Note: Aug = Factor-Augmented, Ind = Unaugmented Individual, RW = Random Walk.
(b) Out-of-Sample UoW and UoW-based DM Test for 7-day Crypto
Asset Model Random Walk AR5 HAR MIDAS (k=30) MIDAS (k=50) MIDAS (k=80) LSTM
Panel A: Out-of-Sample Utility of Wealth (in %)
BTC Individual 2.66 3.31 2.87 3.25 3.23 3.22 3.66
Panel - 3.32 3.22 3.27 3.27 3.27 3.65
Factor-Augmented - 3.35 3.10 3.34 3.33 3.34 3.68
ETH Individual 3.13 3.51 3.46 3.49 3.48 3.48 3.58
Panel - 3.50 3.46 3.48 3.48 3.48 3.60
Factor-Augmented - 3.52 3.47 3.53 3.52 3.51 3.65
XRP Individual 2.72 3.14 3.11 3.08 3.06 3.04 2.36
Panel - 3.10 3.13 3.03 3.03 3.03 2.48
Factor-Augmented - 3.24 3.23 3.20 3.20 3.19 2.41
ADA Individual 3.39 3.62 3.57 3.59 3.59 3.59 3.31
Panel - 3.60 3.61 3.60 3.60 3.60 3.45
Factor-Augmented - 3.65 3.62 3.62 3.62 3.61 3.35
LTC Individual 3.23 3.59 3.52 3.56 3.55 3.55 3.68
Panel - 3.56 3.56 3.56 3.56 3.56 3.69
Factor-Augmented - 3.60 3.57 3.59 3.58 3.58 3.71
Panel B: UoW-based Diebold-Mariano Test Statistics
BTC Aug vs. RW NA 3.35 3.10 3.34 3.33 3.34 3.68
Aug vs. Ind 1.46 1.43 1.96 2.17 2.06 2.93
ETH Aug vs. RW NA 3.52 3.47 3.53 3.52 3.51 3.65
Aug vs. Ind 1.55 0.668 1.84 1.79 1.74 2.38
XRP Aug vs. RW NA 3.24 3.23 3.20 3.20 3.19 2.41
Aug vs. Ind 1.82 1.20 1.52 1.62 1.70 1.63
ADA Aug vs. RW NA 3.65 3.62 3.62 3.62 3.61 3.35
Aug vs. Ind 2.43 1.13 2.17 2.30 2.24 2.32
LTC Aug vs. RW NA 3.60 3.57 3.59 3.58 3.58 3.71
Aug vs. Ind 1.67 1.72 1.89 2.24 1.89 2.44
Note: Definitions same as table (a).

### v3

@@
 Time-Varying Factor-Augmented Models for Volatility Forecasting
-Abstract
+Abstract.
 Accurate volatility forecasts are vital in modern finance for risk management, portfolio allocation, and strategic decision-making.
@@
 Notably, a practical pairs-trading strategy built on our forecasts delivers superior risk-adjusted returns and profitability, particularly under adverse market conditions.
-Keywords Volatility Forecasting, Factor Models, Long Short-Term Memory (LSTM), HAR, MIDAS, Volatility‑Based Pairs Trading
-1 Introduction
+Volatility Forecasting, Factor Models, Long Short-Term Memory (LSTM), HAR, MIDAS, Volatility‑Based Pairs Trading
+copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singapore, Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore, Singaporedoi: 10.1145/3768292.3770407isbn: 979-8-4007-2220-2/2025/11ccs: Applied computing Operations researchccs: Computing methodologies Machine learning
+1. Introduction
 From tick‑by‑tick hedging to long‑term investment strategy, reliable volatility forecasts are indispensable for guiding financial decisions.
-A vast literature—ranging from linear ARCH engle1982autoregressive , GARCH bollerslev1986generalized , and HAR corsi2009simple to nonlinear neural networks such as LSTMs hochreiter1997long —is typically applied on an asset-by-asset basis, modeling each asset in isolation.
-This univariate approach overlooks market-wide shocks that propagate rapidly across assets hamao1990correlations , often biasing forecasts and understating systemic risk.
-Although multivariate extensions such as BEKK‑GARCH engle1995multivariate , DCC‑GARCH engle2002dynamic , Conditional Autoregressive Wishart golosnoy2012conditional , and deep multivariate LSTM architectures capture shared dynamics, they are rarely used in real‑world settings due to parameter explosion, fragile estimation, and heavy tuning burdens bauwens2006multivariate ; engle2019large .
+A vast literature—ranging from linear ARCH (Engle, 1982), GARCH (Bollerslev, 1986), and HAR (Corsi, 2009) to nonlinear neural networks such as LSTMs (Hochreiter and Schmidhuber, 1997)—is typically applied on an asset-by-asset basis, modeling each asset in isolation.
+This univariate approach overlooks market-wide shocks that propagate rapidly across assets (Hamao et al., 1990), often biasing forecasts and understating systemic risk.
+Although multivariate extensions such as BEKK‑GARCH (Engle and Kroner, 1995), DCC‑GARCH (Engle, 2002), Conditional Autoregressive Wishart (Golosnoy et al., 2012), and deep multivariate LSTM architectures capture shared dynamics, they are rarely used in real‑world settings due to parameter explosion, fragile estimation, and heavy tuning burdens (Bauwens et al., 2006; Engle et al., 2019).
 Factor models offer a natural remedy by efficiently summarizing co-movements through a low-dimensional set of latent drivers.
@@
 First, most factor models are designed for returns, not volatility.
-Canonical linear frameworks such as Fama-French and Carhart fama1993common ; carhart1997persistence , and their modern machine‑learning successors gu2020empirical , all distill factors from returns data.
+Canonical linear frameworks such as Fama-French and Carhart (Fama and French, 1993; Carhart, 1997), and their modern machine‑learning successors (Gu et al., 2020), all distill factors from returns data.
 Consequently, the rich patterns of commonality in volatility—which often exhibit their own distinct dynamics—remain largely unexplored.
-Second, although a few recent pioneering studies do extend factor models to realized volatility atak2013factor ; ding2025multiplicative , they still impose static factor loadings, which fail to capture time-varying interdependencies—an especially critical flaw given that volatility co-movement evolves over time del2008dynamic .
-Although recent econometric advances do provide time-varying loadings estimators mikkelsen2019consistent , these techniques have so far been confined exclusively to asset returns.
+Second, although a few recent pioneering studies do extend factor models to realized volatility (Atak and Kapetanios, 2013; Ding et al., 2025), they still impose static factor loadings, which fail to capture time-varying interdependencies—an especially critical flaw given that volatility co-movement evolves over time (Del Negro and Otrok, 2008).
+Although recent econometric advances do provide time-varying loadings estimators (Mikkelsen et al., 2019), these techniques have so far been confined exclusively to asset returns.
 To bridge these gaps, we propose a model-agnostic Factor-Augmented Volatility Forecast framework that captures evolving market-wide volatility shocks with computational tractability.
@@
   • We provide a practical, end-to-end guide to factor extraction, selection, and integration, offering a clear roadmap for tailoring the framework to any forecasting model, horizon, or asset class.
-2 Related Work
+2. Related Work
 Research in discrete-time volatility forecasting has evolved significantly. However, existing models present common trade-offs, which our framework is designed to address.
 Univariate Models.
-Frameworks from early AR models to more sophisticated ARCH engle1982autoregressive , GARCH bollerslev1986generalized ; bollerslev1987conditionally , and HAR corsi2009simple have long been pillars of the field due to their interpretability.
-Extensions have incorporated features like mixed-frequency data conrad2020two ; santos2014volatility and long-memory processes baillie1996long .
-However, by modeling each asset in isolation, these models suffer from a "univariate blind spot," ignoring the systemic co-movements that are crucial drivers of market-wide risk herskovic2016common ; hamao1990correlations .
+Frameworks from early AR models to more sophisticated ARCH (Engle, 1982), GARCH (Bollerslev, 1986, 1987), and HAR (Corsi, 2009) have long been pillars of the field due to their interpretability.
+Extensions have incorporated features like mixed-frequency data (Conrad and Kleen, 2020; Santos and Ziegelmann, 2014) and long-memory processes (Baillie, 1996).
+However, by modeling each asset in isolation, these models suffer from a ”univariate blind spot,” ignoring the systemic co-movements that are crucial drivers of market-wide risk (Herskovic et al., 2016; Hamao et al., 1990).
 Multivariate Models.
 Multivariate models are designed to capture asset interplay directly.
-Within the GARCH family, the BEKK specification engle1995multivariate guarantees positive-definite covariance matrices, while the Dynamic Conditional Correlation (DCC) model engle2002dynamic offers a more feasible estimation process.
-AR-style alternatives like the Conditional Autoregressive Wishart (CAW) golosnoy2012conditional provide further flexibility.
-While theoretically comprehensive, these models are often impractical for large systems due to the explosion of parameter count bauwens2006multivariate and numerically unstable estimation engle2019large .
+Within the GARCH family, the BEKK specification (Engle and Kroner, 1995) guarantees positive-definite covariance matrices, while the Dynamic Conditional Correlation (DCC) model (Engle, 2002) offers a more feasible estimation process.
+AR-style alternatives like the Conditional Autoregressive Wishart (CAW) (Golosnoy et al., 2012) provide further flexibility.
+While theoretically comprehensive, these models are often impractical for large systems due to the explosion of parameter count (Bauwens et al., 2006) and numerically unstable estimation (Engle et al., 2019).
 Machine Learning Models.
 A parallel track has leveraged machine learning to capture the complex, nonlinear volatility patterns.
-Deep learning models like LSTM hochreiter1997long have shown superior performance over classical models by modeling long-range dependencies in financial time series liu2019novel ; mcnally2018predicting .
-However, these models are not immune to the "curse of dimensionality." When applied in a fully multivariate setting, they face excessive parameterization, high overfitting risk, and prohibitive computational costs ge2023review , inheriting the same fundamental limitations as their statistical counterparts.
+Deep learning models like LSTM (Hochreiter and Schmidhuber, 1997) have shown superior performance over classical models by modeling long-range dependencies in financial time series (Liu, 2019; McNally et al., 2018).
+However, these models are not immune to the ”curse of dimensionality.” When applied in a fully multivariate setting, they face excessive parameterization, high overfitting risk, and prohibitive computational costs (Ge et al., 2023), inheriting the same fundamental limitations as their statistical counterparts.
 Factor Models.
 Factor methods are the backbone of modern portfolio theory, distilling hundreds of correlated returns into a handful of common drivers and powering scalable asset allocation, hedging, and attribution.
-Widely used examples range from statistical factor‑GARCH variants such as OGARCH alexander2000primer and GO‑GARCH van2002go to recent factor‑structured deep networks feng2024deep .
-While these models achieve parsimony by using latent factors, most of the existing work targets returns gu2020empirical ; factor models for volatility remain comparatively sparse and always assume time‑invariant loadings atak2013factor ; ding2025multiplicative .
+Widely used examples range from statistical factor‑GARCH variants such as OGARCH (Alexander, 2000) and GO‑GARCH (Van der Weide, 2002) to recent factor‑structured deep networks (Feng et al., 2024).
+While these models achieve parsimony by using latent factors, most of the existing work targets returns (Gu et al., 2020; CHEN et al., 2020; Chen et al., 2019; Chen and Chen, 2022; Liu and Chen, 2022; Chen and Fan, 2023; Chen et al., 2024d, a, c; Kong et al., 2024); factor models for volatility remain comparatively sparse and always assume time‑invariant loadings (Atak and Kapetanios, 2013; Ding et al., 2025).
 Our research directly confronts these trade-offs by proposing a factor augmentation framework.
 This approach enhances existing volatility models with a set of dynamic, cross-sectional factors, enabling the framework to capture systemic, time-varying market interdependence without adding additional computational overhead, thereby balancing predictive accuracy with practical feasibility.
-3 Data and Volatility Measurement
-3.1 Data
+3. Data and Volatility Measurement
+3.1. Data
 The empirical analysis is conducted on two high-frequency datasets, encompassing U.S. equities and major cryptocurrencies.
@@
 Kaiko compiles transaction-level records from major digital asset exchanges, providing comprehensive data including timestamps, trade volumes, and direction.
-3.2 Volatility Measurement
-3.2.1 Daily Realized Volatility
-Following andersen2001distribution , we estimate daily realized volatility (RV) nonparametrically from high-frequency midpoint prices.
+3.2. Volatility Measurement
+3.2.1. Daily Realized Volatility
+Following (Andersen et al., 2001), we estimate daily realized volatility (RV) nonparametrically from high-frequency midpoint prices.
 After filtering out non-positive, crossed, and spurious quotes, midpoint prices for asset $i$ on day $t$ are sampled every five minutes:
-$\text{MID}_{i,t}=\frac{\text{BID}_{i,t}+\text{ASK}_{i,t}}{2}.$ (1)
+(1) $\text{MID}_{i,t}=\frac{\text{BID}_{i,t}+\text{ASK}_{i,t}}{2}.$
 The daily RV is then calculated as the square root of the sum of squared intraday returns:
-$RV_{i,t}=\sqrt{\sum_{s=1}^{N}\left[\ln\left(\frac{\text{MID}_{i,t_{s}}}{\text{MID}_{i,t_{s-1}}}\right)\right]^{2}}\;,$ (2)
+(2) $RV_{i,t}=\sqrt{\sum_{s=1}^{N}\left[\ln\left(\frac{\text{MID}_{i,t_{s}}}{\text{MID}_{i,t_{s-1}}}\right)\right]^{2}}\;,$
 where $\mathrm{MID}_{i,t_{s}}$ is the mid‐price at the $s$th five‐minute interval, $N=78$ for a standard equity trading day and $N=288$ for 24‐hour cryptocurrency markets.
-3.2.2 Aggregated Realized Volatility
-To capture medium-term risk dynamics, we construct weekly realized volatility by averaging daily RVs over a 7-day window ($h=7$) following (corsi2009simple, ; bollerslev2018risk, ):
-$RV_{t}^{7}=\frac{1}{7}\sum_{j=0}^{6}RV_{t-j}\;.$ (3)
+3.2.2. Aggregated Realized Volatility
+To capture medium-term risk dynamics, we construct weekly realized volatility by averaging daily RVs over a 7-day window ($h=7$) following (Corsi, 2009; Bollerslev et al., 2018):
+(3) $RV_{t}^{7}=\frac{1}{7}\sum_{j=0}^{6}RV_{t-j}\;.$
 This 7‑day moving average smooths daily noise, yielding a stable volatility gauge that matches common weekly rebalancing and reporting cycles.
-4 Methodology
-4.1 Volatility Factor Analysis Across Assets
-4.1.1 Time-Varying Factor Model
+4. Methodology
+4.1. Volatility Factor Analysis Across Assets
+4.1.1. Time-Varying Factor Model
 Let $\mathbf{y}_{t}=(RV_{1,t},\dots,RV_{p,t})$ denote the $p$-dimensional vector of realized volatilities at time $t=1,\ldots,T$.
-Following the time‑varying matrix factor model of chen2024timevaryingmatrixfactormodels , we posit the locally‑smooth factor representation:
-$\mathbf{y}_{t}\;=\;\mathbf{\Lambda}_{t}\mathbf{f}_{t}+\mathbf{\varepsilon}_{t},$ (4)
+Following the time‑varying matrix factor model of (Chen et al., 2024b), we posit the locally‑smooth factor representation:
+(4) $\mathbf{y}_{t}\;=\;\mathbf{\Lambda}_{t}\mathbf{f}_{t}+\mathbf{\varepsilon}_{t},$
 where $\mathbf{f}_{t}\in\mathbb{R}^{k}$ is the $k$-dimensional latent factors, and $\mathbf{\varepsilon}_{t}\in\mathbb{R}^{p}$ is the idiosyncratic errors, assumed to be uncorrelated with $\mathbf{f}_{t}$ but allowed to exhibit weak serial dependence.
@@
 Let $\mathbf{\Sigma}_{f}=\mathbb{E}(\mathbf{f}_{t}\mathbf{f}_{t}^{\top})$ and $\mathbf{\Sigma}_{\varepsilon}=\mathbb{E}(\varepsilon_{t}\varepsilon_{t}^{\top})$. The second uncentered moment satisfies
-$\mathbf{\Sigma}_{y,t}:=\mathbb{E}(\mathbf{y}_{t}\mathbf{y}_{t}^{\top})=\mathbf{\Lambda}_{t}\mathbf{\Sigma}_{f}\mathbf{\Lambda}_{t}^{\top}+\mathbf{\Sigma}_{\varepsilon}.$ (5)
-Under mild regularity conditions, the rank‑$k$ signal component $\mathbf{\Lambda}_{t}\mathbf{\Sigma}_{f}\mathbf{\Lambda}_{t}^{\top}$ implies that the top $k$ eigenvectors of $\mathbf{\Sigma}_{y,t}$ span the column space of $\mathbf{\Lambda}_{t}$ up to rotation, as established in chen2024timevaryingmatrixfactormodels .
-In principle, one could estimate $\mathbf{\Sigma}_{y,t}$ using kernel smoothing in $t/T$ as in chen2024timevaryingmatrixfactormodels .
+(5) $\mathbf{\Sigma}_{y,t}:=\mathbb{E}(\mathbf{y}_{t}\mathbf{y}_{t}^{\top})=\mathbf{\Lambda}_{t}\mathbf{\Sigma}_{f}\mathbf{\Lambda}_{t}^{\top}+\mathbf{\Sigma}_{\varepsilon}.$
+Under mild regularity conditions, the rank‑$k$ signal component $\mathbf{\Lambda}_{t}\mathbf{\Sigma}_{f}\mathbf{\Lambda}_{t}^{\top}$ implies that the top $k$ eigenvectors of $\mathbf{\Sigma}_{y,t}$ span the column space of $\mathbf{\Lambda}_{t}$ up to rotation, as established in (Chen et al., 2024b).
+In principle, one could estimate $\mathbf{\Sigma}_{y,t}$ using kernel smoothing in $t/T$ as in (Chen et al., 2024b).
 For practical implementation, we instead adopt a rolling‑window (uniform‑kernel) approximation.
@@
 We apply this time-varying factor model separately to equities and cryptocurrencies, producing two distinct sets of $\widehat{\mathbf{\Lambda}}_{t}\text{ and }\widehat{\mathbf{f}}_{t}$.
-4.1.2 Divergent Factor Structures: Concentration in Crypto vs. Diffusion in Equity
+4.1.2. Divergent Factor Structures: Concentration in Crypto vs. Diffusion in Equity
 Both the equity and cryptocurrency markets exhibit hierarchical factor structures, yet they diverge significantly in risk concentration.
@@
 A single, dominant systemic factor $\hat{\mathbf{f}}_{t}^{(1)}$ serves as the primary market-wide driver, accounting for a commanding 79.1% of total variance.
-This factor is closely linked to the Bitcoin cycle, with its overwhelming influence confirmed by the sharp "elbow" in Figure 1(d).
+This factor is closely linked to the Bitcoin cycle, with its overwhelming influence confirmed by the sharp ”elbow” in Figure 1(d).
 While subsequent factors isolate the dynamics of major altcoins like Ethereum and Ripple, their marginal contributions are more modest, highlighting a market driven by one principal force.
@@
 For short-term (1-day) forecasts, we select only the single, dominant factor for each asset class to maximize the signal-to-noise ratio.
-For medium-term (7-day) forecasts, we employ an explained-variance threshold jolliffe2016principal ; pukthuanthong2009global .
+For medium-term (7-day) forecasts, we employ an explained-variance threshold (Jolliffe and Cadima, 2016; Pukthuanthong and Roll, 2009).
 This involves selecting the minimum number of factors required to meet a predetermined cumulative variance target, which we tailor to each market’s unique structure (e.g., 85% for equities and 90% for crypto).
@@
 (d) Explained Variance for Crypto
-Figure 1: Factor Loadings and Scree Plots for Equities and Crypto
-4.2 Factor-Augmented Forecasting Models
+Figure 1. Factor Loadings and Scree Plots for Equities and Crypto
+4.2. Factor-Augmented Forecasting Models
 To demonstrate the flexibility and model-agnostic nature of our framework, we apply it uniformly to four benchmark models.
@@
 The forecasts are based on two estimation approaches: an expanding window for the three statistical models and a fixed 80:20 train-test split for the more computationally intensive LSTM.
-4.2.1 Autoregressive (AR)
+4.2.1. Autoregressive (AR)
 The AR model regress the target volatility $Y_{t+h}$ on the five most recent daily RVs to capture short-memory dynamics over a full trading week without over-parameterization:
-$Y_{t+h}=\beta_{0}+\sum_{i=0}^{4}\beta_{i}RV_{t-i}+\epsilon_{t}.$ (6)
+(6) $Y_{t+h}=\beta_{0}+\sum_{i=0}^{4}\beta_{i}RV_{t-i}+\epsilon_{t}.$
 Given the AR model’s simple linear structure, the factor-augmented version extends the baseline by directly incorporating the top-$S$ factors as additional predictors:
-$Y_{t+h}\;=\;\beta_{0}+\sum_{i=0}^{4}\beta_{i}RV_{t-i}\;+\;\sum_{j\in[S]}\gamma_{j}\,\widehat{\mathbf{f}}_{j,t}\;+\;\epsilon_{t}.$ (7)
-4.2.2 Heterogeneous Autoregressive (HAR)
-The HAR model captures multi-scale volatility persistence using daily, weekly, and monthly components corsi2009simple :
-$Y_{t+h}=\beta_{0}+\beta^{d}RV_{t}+\beta^{w}RV^{7}_{t}+\beta^{m}RV^{30}_{t}+\epsilon_{t},$ (8)
+(7) $Y_{t+h}\;=\;\beta_{0}+\sum_{i=0}^{4}\beta_{i}RV_{t-i}\;+\;\sum_{j\in[S]}\gamma_{j}\,\widehat{\mathbf{f}}_{j,t}\;+\;\epsilon_{t}.$
+4.2.2. Heterogeneous Autoregressive (HAR)
+The HAR model captures multi-scale volatility persistence using daily, weekly, and monthly components (Corsi, 2009):
+(8) $Y_{t+h}=\beta_{0}+\beta^{d}RV_{t}+\beta^{w}RV^{7}_{t}+\beta^{m}RV^{30}_{t}+\epsilon_{t},$
 where $RV^{d}_{t}$, $RV^{w}_{t}$, and $RV^{m}_{t}$ represent daily, weekly, and monthly volatilities, respectively.
 To align with the HAR’s multi-scale structure, our augmentation introduces factors constructed over corresponding daily and weekly horizons:
-$\begin{split}Y_{t+h}\;=&\;\beta_{0}+\beta^{d}RV_{t}+\beta^{w}RV^{7}_{t}+\beta^{m}RV^{30}_{t}\\ &\;+\;\sum_{j\in[{S}^{d}]}\gamma^{d}_{j}\,\widehat{\mathbf{f}}^{\,d}_{j,t}\;+\;\sum_{r\in[{S}^{w}]}\gamma^{w}_{r}\,\widehat{\mathbf{f}}^{\,w}_{r,t}\;+\;\epsilon_{t},\end{split}$ (9)
+(9) $\begin{split}Y_{t+h}\;=&\;\beta_{0}+\beta^{d}RV_{t}+\beta^{w}RV^{7}_{t}+\beta^{m}RV^{30}_{t}\\ &\;+\;\sum_{j\in[{S}^{d}]}\gamma^{d}_{j}\,\widehat{\mathbf{f}}^{\,d}_{j,t}\;+\;\sum_{r\in[{S}^{w}]}\gamma^{w}_{r}\,\widehat{\mathbf{f}}^{\,w}_{r,t}\;+\;\epsilon_{t},\end{split}$
 where $\widehat{\mathbf{f}}^{d}_{j,t}$ and $\widehat{\mathbf{f}}^{w}_{r,t}$ denote the top-$S^{d}$ daily factors and the top-$S^{w}$ weekly factors, respectively.
-4.2.3 Mixed Data Sampling (MIDAS)
-The MIDAS model captures long-memory dynamics by incorporating many past observations through a parsimonious weighting scheme, specified as a Beta-polynomial distributed lag ghysels2007midas :
-$Y_{t+h}=\beta_{0}+\beta_{1}\left[a(1)^{-1}a(L)RV_{t}\right]+\epsilon_{t},$ (10)
+4.2.3. Mixed Data Sampling (MIDAS)
+The MIDAS model captures long-memory dynamics by incorporating many past observations through a parsimonious weighting scheme, specified as a Beta-polynomial distributed lag (Ghysels et al., 2007):
+(10) $Y_{t+h}=\beta_{0}+\beta_{1}\left[a(1)^{-1}a(L)RV_{t}\right]+\epsilon_{t},$
 where lag weights $a(L)$ follow:
-$a_{i}=\left(\frac{i}{k}\right)^{\theta_{1}-1}\left(1-\frac{i}{k}\right)^{\theta_{2}-1}\frac{\Gamma(\theta_{1}+\theta_{2})}{\Gamma(\theta_{1})\Gamma(\theta_{2})},\quad i=1,\ldots,k,$ (11)
+(11) $a_{i}=\left(\frac{i}{k}\right)^{\theta_{1}-1}\left(1-\frac{i}{k}\right)^{\theta_{2}-1}\frac{\Gamma(\theta_{1}+\theta_{2})}{\Gamma(\theta_{1})\Gamma(\theta_{2})},\quad i=1,\ldots,k,$
 with normalization $a(1)=\sum_{i=1}^{k}a_{i}$. $\theta_{1}=1$ and $\theta_{2}$ are optimized via in-sample grid search. Default $k=30$ is extended with $k\in\{30,50,80\}$ for longer horizons.
 To align with the MIDAS’s core principle of parsimoniously weighting long-lag structures, we augment each factor through its own analogous Beta-polynomial $b(L)$:
-$Y_{t+h}\;=\;\beta_{0}+\beta_{1}\!\bigl[a(1)^{-1}a(L)RV_{t}\bigr]+\sum_{j\in[S]}\gamma_{j}\bigl[b(1)^{-1}b(L)\,\widehat{\mathbf{f}}_{j,t}\bigr]+\epsilon_{t}.$ (12)
-4.2.4 Long Short-Term Memory (LSTM)
+(12) $Y_{t+h}\;=\;\beta_{0}+\beta_{1}\!\bigl[a(1)^{-1}a(L)RV_{t}\bigr]+\sum_{j\in[S]}\gamma_{j}\bigl[b(1)^{-1}b(L)\,\widehat{\mathbf{f}}_{j,t}\bigr]+\epsilon_{t}.$
+4.2.4. Long Short-Term Memory (LSTM)
 Long Short‑Term Memory network is a gated recurrent architecture designed to capture long‑range dependencies in time‑series data.
@@
 The final hidden state feeds a dense output layer that forecasts the target volatility $Y_{t+h}$.
-4.3 Evaluation Metrics
+4.3. Evaluation Metrics
 Model performance is examined through three lenses: (i) statistical accuracy, (ii) statistical significance, and (iii) economic value.
-4.3.1 Prediction Accuracy
-Prediction accuracy is assessed with three widely used metrics: $R^{2}$, MSE, and QLIKE, which respectively capture explained variance, average prediction error, and distributional accuracy respectively patton2011volatility ; zhang2024volatility :
-$R^{2}$ $=1-\sum_{t}\left(\frac{RV^{h}_{t+h}-\widehat{RV}^{h}_{t+h}}{RV^{h}_{t+h}-\overline{RV}^{h}_{t+h}}\right)^{2},$ (13)
-$\mathrm{MSE}$ $=\frac{1}{T}\sum_{t}\left(RV^{h}_{t+h}-\widehat{RV}^{h}_{t+h}\right)^{2},$ (14)
-$\mathrm{\mathrm{QLIKE}}$ $=\frac{1}{T}\sum_{t}\!\left[\frac{RV^{h}_{t+h}}{\widehat{RV}^{h}_{t+h}}-\ln\!\left(\frac{RV^{h}_{t+h}}{\widehat{RV}^{h}_{t+h}}\right)-1\right],$ (15)
-4.3.2 Economic Implication
+4.3.1. Prediction Accuracy
+Prediction accuracy is assessed with three widely used metrics: $R^{2}$, MSE, and QLIKE, which respectively capture explained variance, average prediction error, and distributional accuracy respectively (Patton, 2011; Zhang et al., 2024):
+(13) $R^{2}$ $=1-\sum_{t}\left(\frac{RV^{h}_{t+h}-\widehat{RV}^{h}_{t+h}}{RV^{h}_{t+h}-\overline{RV}^{h}_{t+h}}\right)^{2},$
+(14) $\mathrm{MSE}$ $=\frac{1}{T}\sum_{t}\left(RV^{h}_{t+h}-\widehat{RV}^{h}_{t+h}\right)^{2},$
+(15) $\mathrm{\mathrm{QLIKE}}$ $=\frac{1}{T}\sum_{t}\!\left[\frac{RV^{h}_{t+h}}{\widehat{RV}^{h}_{t+h}}-\ln\!\left(\frac{RV^{h}_{t+h}}{\widehat{RV}^{h}_{t+h}}\right)-1\right],$
+4.3.2. Economic Implication
 We assess practical relevance using the out-of-sample Utility-of-Wealth ($\mathrm{UoW}$) metric:
-$\widehat{\mathrm{UoW}}=\frac{1}{T}\sum_{t=1}^{T}\!\left[\frac{SR^{2}}{\gamma}\frac{RV^{7}_{t+7}}{\widehat{RV}^{7}_{t+7}}-\frac{SR^{2}}{2\gamma}\!\left(\frac{RV^{7}_{t+7}}{\widehat{RV}^{7}_{t+7}}\right)^{2}\right],$ (16)
-where we follow bollerslev2018risk in setting the target Sharpe ratio to $\mathrm{SR}=0.4$, a typical long-run estimate for volatility-timing strategies, and choose risk aversion $\gamma=2$ to represent moderate investor preferences.
+(16) $\widehat{\mathrm{UoW}}=\frac{1}{T}\sum_{t=1}^{T}\!\left[\frac{SR^{2}}{\gamma}\frac{RV^{7}_{t+7}}{\widehat{RV}^{7}_{t+7}}-\frac{SR^{2}}{2\gamma}\!\left(\frac{RV^{7}_{t+7}}{\widehat{RV}^{7}_{t+7}}\right)^{2}\right],$
+where we follow (Bollerslev et al., 2018) in setting the target Sharpe ratio to $\mathrm{SR}=0.4$, a typical long-run estimate for volatility-timing strategies, and choose risk aversion $\gamma=2$ to represent moderate investor preferences.
 Higher $\mathrm{UoW}$ denotes greater expected utility for a volatility-timing investor.
-4.3.3 Statistical Significance
-We assess systematic improvements with the Diebold–Mariano testdiebold2002comparing . For each horizon $h$, we compute the one‐period loss differential:
-$d_{t}\;=\;L\bigl(RV^{h}_{t+h},\widehat{RV}^{h}_{t+h,\mathrm{aug}}\bigr)\;-\;L\bigl(RV^{h}_{t+h},\widehat{RV}^{h}_{t+h,\mathrm{RW}}\bigr),$ (17)
+4.3.3. Statistical Significance
+We assess systematic improvements with the Diebold–Mariano test(Diebold and Mariano, 2002). For each horizon $h$, we compute the one‐period loss differential:
+(17) $d_{t}\;=\;L\bigl(RV^{h}_{t+h},\widehat{RV}^{h}_{t+h,\mathrm{aug}}\bigr)\;-\;L\bigl(RV^{h}_{t+h},\widehat{RV}^{h}_{t+h,\mathrm{Baseline}}\bigr),$
 where
@@
 The $\mathrm{DM}$ statistic:
-$\mathrm{DM}=\frac{\bar{d}}{\sqrt{\widehat{\mathrm{Var}}(\bar{d})}},\quad\bar{d}=\frac{1}{T}\sum_{t=1}^{T}d_{t},$ (19)
+(19) $\mathrm{DM}=\frac{\bar{d}}{\sqrt{\widehat{\mathrm{Var}}(\bar{d})}},\quad\bar{d}=\frac{1}{T}\sum_{t=1}^{T}d_{t},$
 is asymptotically $\mathcal{N}(0,1)$ under $H_{0}:E[d_{t}]=0$. A positive DM value indicates that the factor-augmented forecasts significantly outperform the unaugmented benchmarks.
-5 Experiment Result
-Table 1: Cryptocurrency Volatility Forecast Accuracy: 1- and 7-Day Out-of-Sample Results
+5. Experiment Result
+Table 1. Cryptocurrency Volatility Forecast Accuracy: 1- and 7-Day Out-of-Sample Results
 Panel A: Crypto in 1‐Day Horizon
@@
 LSTM‐Aug 61.66 1.30 0.0374 51.89 2.03 0.0385 44.25 1.41 0.1186 41.16 6.19 0.0616 59.83 2.38 0.0340
-Note: All MSE values are in units of $10^{-4}$.   Improvement:  $\geq+15\%$  $\geq+10\%$  $\geq+5\%$  $0\%$  $\leq-5\%$
-5.1 A Universal Performance Boost Across All Models and Horizons
+Note: All MSE values are reported in units of $10^{-4}$.  Improvement:   $\geq+15\%$     $\geq+10\%$    $\geq+5\%$     > $0\%$     $\leq-5\%$
+5.1. A Universal Performance Boost Across All Models and Horizons
 Factor augmentation universally and significantly boosts forecast accuracy across every model, asset class, and horizon tested, proving its effectiveness for both short- and medium-term predictions, as shown in Tables 1 and 4.
@@
 These results demonstrate the versatile power of factor augmentation, which consistently improves forecast accuracy, especially when a model’s architecture aligns with the task horizon.
-5.2 Value Across Markets: Robust in Equities, Super‑Charged in Crypto
+5.2. Value Across Markets: Robust in Equities, Super‑Charged in Crypto
 Factor augmentation framework proves its power and versatility across fundamentally different market structures, delivering robust gains in the diffuse equity market and a super-charged performance boost in the highly concentrated crypto market.
@@
 These results confirm that factor augmentation is not only a reliable enhancer in different markets, but also becomes especially powerful when risk drivers are concentrated.
-5.3 Statistically and Economically Significant Gains
+5.3. Statistically and Economically Significant Gains
 Factor augmentation delivers gains that are both substantial and statistically significant in both predictive accuracy and economic value.
 We demonstrate this on the crypto 7-day horizon task, where our framework shows the largest improvements.
-Three unaugmented baselines are employed: the Random Walk, an "Individual" model with asset-specific coefficients, and an "Panel" model that share coefficients across assets.
+Three unaugmented baselines are employed: the Random Walk, an ”Individual” model with asset-specific coefficients, and an ”Panel” model that share coefficients across assets.
 The framework’s edge in predictive accuracy is both substantial and statistically unequivocal.
@@
 As detailed in Table 5(b), the augmented models consistently deliver unambiguously higher UoW outcomes, and a corresponding UoW-based DM test formally validates that these gains are statistically significant, providing clear evidence of the framework’s real-world benefits for investors.
-6 Portfolio Backtest
+6. Portfolio Backtest
 To rigorously assess the framework’s financial value, we conduct a volatility‑scaled pair‑trading backtest.
-6.1 Cointegration Test
+6.1. Cointegration Test
 To identify a stationary, mean-reverting spread suitable for our pairs-trading strategy, we perform a two-stage cointegration analysis on the price series within each market.
 We apply the Johansen trace test to each five-asset group. While it finds no evidence of cointegration at any rank for the equity group, the crypto sector indicates a cointegration rank of $r=2$.
-Table 2: Cointegrating Vector Coefficients for Crypto
+Table 2. Cointegrating Vector Coefficients for Crypto
 ADA BTC ETH LTC XRP
@@
 Consequently, our backtest focuses exclusively on the ADA–ETH spread, excluding all other pairs.
-6.2 Volatility‑Based Pairs Trading
+6.2. Volatility‑Based Pairs Trading
 Our backtest framework simulates a pairs trading strategy on the cointegrated spread between ADA-ETH. The log-price spread $S_{t}$ is defined as:
-$S_{t}=\ln P_{\text{ADA},t}-\beta\,\ln P_{\text{ETH},t},$ (20)
+(20) $S_{t}=\ln P_{\text{ADA},t}-\beta\,\ln P_{\text{ETH},t},$
 where the hedge ratio $\beta$ is estimated via rolling OLS regression. The performance is evaluated for all four benchmark models and is measured by annualized return and Sharpe ratio.
 The performance is evaluated by final equity, cumulative return, and annualized Sharpe ratio.
-6.2.1 Trading Signal Generation
+6.2.1. Trading Signal Generation
 Trading signals are generated from a rolling z-score of the spread, where $Z_{t}=\frac{S_{t}-\mu_{S_{t}}}{\sigma_{S_{t}}}$.
-We initiate a long position when $Z_{t}$ < -1.5 and a short position when $Z_{t}$ > +1.5, closing the position when the z-score reverts to zero.
+We initiate a long position when $Z_{t}$ ¡ -1.5 and a short position when $Z_{t}$ ¿ +1.5, closing the position when the z-score reverts to zero.
 To limit turnover, each trade is held for at least one day.
 The rolling window length ($W$) is set to 70 days for the statistical models to balance responsiveness to regime changes with statistical stability, while a shorter window of 30 days is used for the LSTM to align with its input structure.
-6.2.2 Volatility-based Position Sizing
+6.2.2. Volatility-based Position Sizing
 Our strategy employs a volatility-targeting approach, where daily position sizes are scaled to achieve a 25% annualized volatility target.
 Each day, we estimate the spread volatility using our model’s forecasts for the component assets’ volatility $\widehat{\sigma}_{\text{ADA},t}$, $\widehat{\sigma}_{\text{ETH},t}$ and their return covariance:
-$\widehat{\sigma}_{\text{spread},t}\;=\;\sqrt{252}\cdot\sqrt{\widehat{\sigma}_{\text{ADA},t}^{2}+\beta^{2}\widehat{\sigma}_{\mathrm{ETH},t}^{2}-2\beta\,\widehat{\mathrm{Cov}}(r_{\text{ADA},t},r_{\text{ETH},t})}.$ (21)
+(21) $\widehat{\sigma}_{\text{spread},t}\;=\;\sqrt{252}\cdot\sqrt{\widehat{\sigma}_{\text{ADA},t}^{2}+\beta^{2}\widehat{\sigma}_{\mathrm{ETH},t}^{2}-2\beta\,\widehat{\mathrm{Cov}}(r_{\text{ADA},t},r_{\text{ETH},t})}.$
 The position size is then set based on this forecast, subject to a 5:1 maximum leverage.
 The backtest is initialized with an equity of \$50,000 and incorporates a 5 basis point transaction cost per leg.
-6.3 Backtest Results
+6.3. Backtest Results
 To ensure a fair comparison, we evaluate statistical models on the full sample with an expanding window and the LSTM on a fixed 80:20 split.
 Each approach is benchmarked against a corresponding Random Walk baseline, with full results presented in Table 3.
-Table 3: Out-of-Sample Backtest for ADA-ETH Pair
+Table 3. Out-of-Sample Backtest for ADA-ETH Pair
 Panel A: Expanding Window Models
@@
 Ann. Sharpe Ratio -0.467 0.787 268.52%
-6.3.1 Dual Gains: Amplifying Return and Sharpening Risk
+6.3.1. Dual Gains: Amplifying Return and Sharpening Risk
 Factor augmentation decisively boosts both profitability and risk-adjusted performance across all statistical models we examine.
@@
 It consistently amplifies reward, sharpens risk-adjusted performance, and sets a new benchmark for the field.
-6.3.2 Robustness in All Market Conditions: Enhancing Gains, Reversing Losses.
+6.3.2. Robustness in All Market Conditions: Enhancing Gains, Reversing Losses.
 Our framework proves its value in all market conditions, systematically enhancing gains in favorable periods while acting as a powerful shield during downturns.
@@
 This remarkable dual capability—to amplify gains in good times and decisively reverse losses in bad—undeniably proves our framework’s practical utility for navigating the complexities of any market conditions.
-7 Conclusion
+7. Conclusion
 This paper proposes a novel, model-agnostic Factor-Augmented Volatility Forecast framework, resolving the trade-off between computationally infeasible multivariate models, overly simplistic univariate models, and rigid static-factor models.
-By extracting a compact set of dynamic factors and their time-varying loadings directly from realized volatilities, our framework allows any baseline model to efficiently capture evolving market-wide co-movements.
+By extracting a compact set of dynamic factors and their time-varying loadings directly from realized volatilities, our frasmework allows any baseline model to efficiently capture evolving market-wide co-movements.
 The framework’s robust performance—evidenced by substantial gains in forecast accuracy and superior risk-adjusted returns—combined with its computational efficiency and interpretability, establishes it as a valuable and practical solution for dynamic risk management.
-Appendix
-8 Equity Market: Forecast Performance
-Table 4: Equity Volatility Forecast Accuracy: 1- and 7-Day Out-of-Sample Results
+Acknowledgements.
+Elynn Chen’s research is supported in part by the NSF Award 2412577.
+Appendix A Equity Market: Forecast Performance
+Table 4. Equity Volatility Forecast Accuracy: 1- and 7-Day Out-of-Sample Results
 Panel A: Equity in 1‐Day Horizon
@@
 LSTM‐Aug 47.36 3.51 0.0964 44.82 6.91 0.0326 45.15 14.54 0.0766 42.74 33.29 0.1416 50.55 12.74 0.0628
-Note: All MSE values are in units of $10^{-4}$.   Improvement:  $\geq+15\%$  $\geq+10\%$  $\geq+5\%$  $0\%$  $\leq-5\%$
-9 Statistical Significance
-(a) Out-of-Sample $R^{2}$ and MSE-based DM Test for 7-day Crypto
+Note: All MSE values are reported in units of $10^{-4}$.  Improvement:   $\geq+15\%$     $\geq+10\%$    $\geq+5\%$     > $0\%$     $\leq-5\%$
+Appendix B Statistical Significance
+(a) Out-of-Sample $R^{2}$ and MSE-based DM Test for 7-day Crypto.
 Asset Model Random Walk AR5 HAR MIDAS (k=30) MIDAS (k=50) MIDAS (k=80) LSTM